OpenStudy (anonymous):

simplify (x – 2) (3x-4)

OpenStudy (anonymous):

3x^2 - 10 + 6

OpenStudy (anonymous):

thank you foe helping me..again :)

OpenStudy (anonymous):

for**

OpenStudy (anonymous):

np

OpenStudy (unklerhaukus):

(x – 2) (3x-4) = x (3x -4) - 2 (3x-4) = 3x^2 -4x - 6x + 8 = 3x^2 -10x + 8